# A SIMPLE AERIAL DETECTION BASELINE OF MULTIMODAL LANGUAGE MODELS

### v1

@@ I Introduction
Recently, MLMs have been introduced into the RS field. While existing RS MLMs have emphasized visual localization capabilities, their detection performance still requires significant improvement. Geochat [2] presents the first open-source RS MLLM learned on tasks including scene classification, image captioning, region description, visual question answering, and visual grounding. The proposed model demonstrates the ability to detect partial objects corresponding to textual instructions but is unable to perform more intensive detection. EarthGPT [3] extends visual inputs with multisensor RS modalities covering optical, synthetic aperture radar, infrared data. While they achieve detection results comparable to traditional detectors, their model is limited to single-class detection performance. SkySenseGPT [4] introduces scene graph generation to RS MLMs to enhance ability of understanding relations between objects. They also provide tolerance detection accuracies under a lower threshold of intersection over union (IoU). We consider that RS MLMs still require further investigation in aerial detection to fully explore their potential.
@@
In this paper, we present a simple baseline for aerial detection, with a focus on enhancing the detection capabilities of MLMs in the RS domain for the first time, named LMMRotate. Specifically, we propose a straightforward method to supervised fine-tune MLMs, achieving detection performance comparable to that of conventional detectors. The MLLM detectors produce parsable text outputs, offering both flexibility and expandability as multi-dataset joint-trained detectors. We then observe that the advantage of conventional detectors in detection metrics largely comes from object confidence scores. To address this, we propose an appropriate evaluation scheme that enables fair comparisons between MLMs and conventional detectors. Additionally, we conduct exploratory experiments using open-source MLMs, demonstrating the potential of MLMs as aerial detectors. We sincerely hope that our work contributes to advancing more comprehensive abilities of MLMs in the RS field.
@@ III Experiment
Joint training (concat) 55.5 - 58.2 70.4 26.6 43.5 36.5 42.3 - -
Joint training (balanced) 56.8 - 58.9 71.2 26.4 43.0 29.6 38.5 - -
On five datasets.  On four datasets, currently excluding RSAR (will be updated with latest results soon).

### v2

@@ I Introduction
Recently, MLMs have been introduced into the RS field. While existing RS MLMs have emphasized visual localization capabilities, their detection performance still requires significant improvement. Geochat [2] presents the first open-source RS MLM learned on tasks including scene classification, image captioning, region description, visual question answering, and visual grounding. The proposed model demonstrates the ability to detect partial objects corresponding to textual instructions but is unable to perform more intensive detection. EarthGPT [3] extends visual inputs with multisensor RS modalities covering optical, synthetic aperture radar, infrared data. While they achieve detection results comparable to traditional detectors, their model is limited to single-class detection performance. SkySenseGPT [4] introduces scene graph generation to RS MLMs to enhance ability of understanding relations between objects. They also provide tolerance detection accuracies under a lower threshold of intersection over union (IoU). We consider that RS MLMs still require further investigation in aerial detection to fully explore their potential.
@@
In this paper, we present a simple baseline for aerial detection, with a focus on enhancing the detection capabilities of MLMs in the RS domain for the first time, named LMMRotate. Specifically, we propose a straightforward method to supervised fine-tune MLMs, achieving detection performance comparable to that of conventional detectors. The MLM detectors produce parsable text outputs, offering both flexibility and expandability as multi-dataset joint-trained detectors. We then observe that the advantage of conventional detectors in detection metrics largely comes from object confidence scores. To address this, we propose an appropriate evaluation scheme that enables fair comparisons between MLMs and conventional detectors. Additionally, we conduct exploratory experiments using open-source MLMs, demonstrating the potential of MLMs as aerial detectors. We sincerely hope that our work contributes to advancing more comprehensive abilities of MLMs in the RS field.
@@ III Experiment
Joint training (concat) 44.5 - 48.7 62.9 21.9 38.3 14.4 20.5 59.2 65.0
Joint training (balanced) 56.1 - 56.3 69.5 25.2 41.8 31.4 37.4 60.8 65.8